OpenStudy (anonymous):

A quadrilateral contains two equal sides measuring 12 cm each with an included right angle. If the measure of the third side is 8 cm and the angle opposite the right angle is 120 degrees, find the fourth side and the area of the quadrlateral. how can I answer this?

OpenStudy (anonymous):

let AB=12cm BC=12cm CD=8cm DE=xcm <ABC=90 <CDA=120 by pythagoras theorem AC=sqrt(12^2+12^2)=12sqrt2 cos <CDA=(8^2+x^2-(12sqrt2)^2)/2*8*x cos120=cos(180-60)=-cos60=(64+x^2-288)/16x -1/2*16x=x^2-224 x^2+8x-224=0 x=(-8+-sqrt(64-4*1*-224))/2*1 x=(-8+-sqrt960)/2 x=(-8+2sqrt240)/2 (as xis positive so taking only positive sign) x=-4+4sqrt15=4(sqrt15-1)

OpenStudy (anonymous):

you use cos CDA but i know that it only use if that triangle is right triangle so it can't be :(

OpenStudy (anonymous):

produce CD to E and draw AE perpendicularCE <EDA=180-120=60 sin 60=AE/AD sqrt3/2=AE/4(sqrt15-1) AE=sqt3/2 *4(sqrt15-1)=2(sqrt45-sqrt3) area=1/2*12*12+1/2*8*2(sqrt45-sqrt3) solve it. I have used <ABC for pythagoras theorem.

OpenStudy (anonymous):

i have used <CDA for cosine formula ,it can be used for any triangle.

OpenStudy (anonymous):

my equation and Draw are not working so i can't make a diagram.

OpenStudy (anonymous):

me also hahah! i think that figure is kite

OpenStudy (ikram002p):

u might need more hypotheses !

OpenStudy (anonymous):



OpenStudy (ikram002p):

humm i cant see the sketch XD

OpenStudy (anonymous):